OpenStudy (anonymous):

iWill Give A Medal To ANYONE Who Can Answer This , What is the equation in point-slope form of the line passing through (3, 6) and (–2, 1)?

OpenStudy (anonymous):

y + 1 = –1(x + 2) y – 6 = –1(x + 3) y + 6 = 1(x + 3) y – 6 = 1(x – 3)

OpenStudy (anonymous):

y-6 = 1(x-3)

OpenStudy (anonymous):

Learn the formula. It is really straight forward. y2 - y1 = m(x2 - x1) in your question you have (x1=3, y1=6) and x2=-2, y2=1) 1-6=m(-2-3) y-6 = m(x-3)